### v1

# On Bollobás-type theorems of $d$-tuples

Abstract
In 1965, Bollobás [3] proved that for a Bollobás set-pair system $\{(A_{i},B_{i})\mid i\in[m]\}$, the maximum value of $\sum_{i=1}^{m}\binom{|A_{i}|+|B_{i}|}{A_{i}}^{-1}$ is $1$. In 2023, Hegedüs and Frankl [7] generalize the concept of Bollobás system to $d$-tuples, and made a conjecture: Suppose $\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a Bollobás system of $d$-tuples, the maximum value of $\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}$ is also $1$. In this paper, we show this conjecture is not correct, and we give an upper bound of it. In the case $d=3$, our upper bound is tight in the mean of asymptotic. We also improve an inequality for skew Bollobás systems of $d$-tuples in [7], and determine the maximum size or a uniform skew Bollobás system.
1 Introduction
Let $[n]=\{1,\ldots,n\}$ be a set of cardinality $n$. The following concepts of Bollobás systems and skew Bollobás systems were introduced by Bollobás [3] and Frankl [4] respectively.
Definition 1.1.
Suppose $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ is a family of pairs of sets, where $A_{i},B_{i}\subseteq[n]$, and $A_{i}\cap B_{i}=\emptyset$. Then $\mathcal{P}$ is called a Bollobás system if $A_{i}\cap B_{j}\neq\emptyset,\forall i\neq j$, and a skew Bollobás system if $A_{i}\cap B_{j}\neq\emptyset,\forall i<j$.
To solve a problem on hypergraphs, Bollobás [3] proved the following theorem in 1965. In 1975, Tarján [19] gave a more elegant proof for it independently.
Theorem 1.2 (Bollobás [3]).
Let $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ be a Bollobás system, where $A_{i},B_{i}\subseteq[n]$. Then
$\sum_{i=1}^{m}\binom{|A_{i}|+|B_{i}|}{|A_{i}|}^{-1}\leqslant 1.$
Theorem 1.2 also gives the maximum size of a uniform Bollobás system. Actually, suppose $|A_{i}|=a,|B_{i}|=b$ for every $i\in[m]$. Then theorem 1.2 shows that $m\leqslant\binom{a+b}{a}$. In 1982, using polynomial method, Frankl [4] proved that this upper bound remains true for skew Bollobás systems.
Theorem 1.3 (Frankl [4]).
Let $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ be a skew Bollobás system, where $A_{i},B_{i}\subseteq[n]$, and $|A_{i}|=a,|B_{i}|=b$ for any $i$. Then $m\leqslant\binom{a+b}{a}$.
While theorem 1.2 no longer hold if we replace the condition Bollobás system by skew Bollobás system. The following (tight) inequalities for skew Bollobás systems are proved by Hegedüs and Frankl [7], and Yue [23].
Theorem 1.4 (Hegedüs and Frankl [7], Yue [23]).
Suppose $A_{i},B_{i}\subseteq[n],i\in[m]$, and $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ is a skew Bollobás system. Then we have
$\sum_{i=1}^{m}\left((1+|A_{i}|+|B_{i}|)\binom{|A_{i}|+|B_{i}|}{|A_{i}|}\right)^{-1}\leqslant 1,$
and so
$\sum_{i=1}^{m}\binom{|A_{i}|+|B_{i}|}{|A_{i}|}^{-1}\leqslant 1+n.$
One natural way to generalize the concept of (skew) Bollobás systems is to ask what will happen if we have a family of $d$-tuples instead of set-pairs. In [7], Hegedüs and Fnankl define Bollobás systems and skew Bollobás systems in the following way.
Definition 1.5.
Let =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ be a collection of $d$-tuples. Suppose that $A_{i}^{(p)}\cap A_{i}^{(q)}=\emptyset$ for any $i\in[m]$ and $p\neq q$. Then $\mathcal{F}$ is a Bollobás system (skew Bollobás system) if for any $i\neq j$ ($i<j$), there exist $p<q$ such that $A_{i}^{(p)}\cap A_{j}^{(q)}\neq\emptyset$.
For a $d$-tuple $(A^{(1)},\ldots,A^{(d)})$, we say $(|A^{(1)}|,\ldots,|A^{(d)}|)$ is its type. Following is an example of Bollobás system of $d$-tuples.
Example 1.
For fixed $a_{1},\ldots,a_{d}$, let $\mathcal{F}$ be the collection of all disjoint $d$-tuples $(A^{(1)},\ldots,A^{(d)})$ of type $(a_{1},\ldots,a_{d})$ such that $A^{(k)}\in[a_{1}+\cdots+a_{d}],\forall k\in[d]$. For distinct $(A^{(1)},\ldots,A^{(d)})$ and $(B^{(1)},\ldots,B^{(d)})$, there is a $k$ such that $A^{(k)}\neq B^{(k)}$. Then we have $A^{(k)}\cap\left(\cup_{l\neq k}B^{(l)}\right)\neq\emptyset$ and $B^{(k)}\cap\left(\cup_{l\neq k}A^{(l)}\right)\neq\emptyset$. Hence $\mathcal{F}$ is a Bollobás system of $d$-tuples.
To generalize theorem 1.2, Hegedüs and Frankl [7] made the following conjecture.
Conjecture 1 (Hegedüs and Frankl [7]).
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a Bollobás system of $d$-tuples. Then
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}}^{-1}\leqslant 1.$
However, we can deny this conjecture with the following example.
Example 2.
Fixed an arbitrary $n$, and $d=3$. For $l=0,\dots,\lfloor\frac{n}{2}\rfloor$, let $\mathcal{F}_{l}$ be the collection of all disjoint triples $(A,B,C)$ of type $(l,n-2l,l)$, such that $A,B,C\subseteq[n]$, and $\mathcal{F}$ be the union of $\mathcal{F}_{l}$’s. We claim that $\mathcal{F}$ is a Bollobás system. Actually, for two different families $(A,B,C),(A^{\prime},B^{\prime},C^{\prime})\in\mathcal{F}_{l}$, if $A=A^{\prime}$, then $B\cap C^{\prime}\neq\emptyset$ and $B^{\prime}\cap C\neq\emptyset$. If $A\neq A^{\prime}$, then $A\cap(B^{\prime}\cup C^{\prime})=\neq\emptyset$ and $A^{\prime}\cap(B\cup C)\neq\emptyset$. For $(A,B,C)\in\mathcal{F}_{l}$ and $(A^{\prime},B^{\prime},C^{\prime})\in\mathcal{F}_{l^{\prime}}$ where $l<l^{\prime}$, note that $|A|+|B|+|C^{\prime}|=|A^{\prime}|+|B|+|C|=n-l+l^{\prime}>n$, so $A\cap(B^{\prime}\cup C^{\prime})\neq\emptyset$ and $A^{\prime}\cap(B\cup C)\neq\emptyset$. Hence $\mathcal{F}$ is a Bollobás system. But we have
$\sum_{(A,B,C)\in\mathcal{F}}{\binom{|A|+|B|+|C|}{|A|,|B|,|C|}}^{-1}=\sum_{l=0}^{\lfloor\frac{n}{2}\rfloor}\sum_{(A,B,C)\in\mathcal{F}_{l}}{\binom{|A|+|B|+|C|}{|A|,|B|,|C|}}^{-1}=\left\lfloor\frac{n+1}{2}\right\rfloor.$
Generalizing the probabilistic method in [23], we can prove the following inequalities for Bollobás systems of $d$-tuples. The inequality for $d=3$ shows the construction in example 2 is nearly extremum, and this inequality is tight as $n\rightarrow\infty$. For the case $d>3$, it is still not clear if the inequality is tight or not.
Theorem 1.6.
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)}\mid i\in[m])\}$ is a Bollobás system of $d$-tuples, and $A_{i}^{(p)}\subseteq[n]$ for all $i\in[m],p\in[d]$. Then for the case $d=3$ we have
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}}^{-1}\leqslant\frac{n+3}{2},$
and for arbitrary $d$ we have
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}}^{-1}\leqslant\frac{1}{d-1}\binom{n+d-2}{d-2}+O(n^{d-3})$
as $n\rightarrow\infty$.
For skew Bollobás systems of $d$-tuples, Hegedüs and Frankl [7] proved the following theorem.
Theorem 1.7 (Hegedüs and Frankl [7]).
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a skew Bollobás system of $d$-tuples, and $A_{i}^{(p)}\subseteq[n]$ for all $i\in[m]$ and $p\in[d]$. Then
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}}^{-1}\leqslant\binom{n+d-1}{d-1}.$
Although this result is already tight, using similar probabilistic argument, we can still improve it as follows.
Theorem 1.8.
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a skew Bollobás system of $d$-tuples. Then
$\sum_{i=1}^{m}\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}\leqslant 1.$
Theorem 1.8 also gives an upper bound of uniform skew Bollobás systems. If we further assume that $|A_{i}^{(p)}|=a_{p},\forall i\in[m],p\in[d]$, then we have $m\leqslant\binom{n+d-1}{d-1}\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}$ where $n$ is the size of the ground set. However, this upper bound is not so good. Actually, we can prove the following upper bound, which is confirmed to be tight by example 1. (Note that a Bollobás system is also a skew Bollobás system.)
Theorem 1.9.
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a skew Bollobás system of $d$-tuples, and $|A_{i}^{(p)}|=a_{p},\forall i\in[m],p\in[d]$. Then
$m\leqslant\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}.$
This paper is organized as follows: theorem 1.6 and theorem 1.8 are proved in section 2, and theorem 1.9 is proved in section 3.
2 Proof of theorem 1.6 and theorem 1.8
Proof of theorem 1.8.
Without lose of generality, we assume $A_{i}^{(k)}\subseteq[n],\forall i\in[m],k\in[d]$ for some $n$. Pick a random permutation $\sigma\in S_{n+d-1}$ uniformly, and it gives the $n+d-1$ elements a new order. We will use the elements in $\{n+1,\ldots,n+d-1\}$ as delimiters. Let $E_{i}$ be the event that elements in $A_{i}^{(1)},\ldots,A_{i}^{(d)}$ appear in order, and each pair of neighbor groups are divided by a delimiter. Formally speaking,
$\begin{split}E_{i}=&\{\sigma\in S_{n+d-1}\mid\sigma(a_{1})<\sigma(b_{1})<\sigma(a_{2})<\cdots<\sigma(b_{d-1})<\sigma(a_{n}),\forall a_{k}\in A_{i}^{(k)},\\ &\textrm{where}~{}\{b_{1},\ldots,b_{d-1}\}=\{n+1,\ldots,n+d-1\}\}.\end{split}$
To calculate the probability of $E_{i}$, we only need to consider about elements in $A_{i}^{(1)}\cup\cdots\cup A_{i}^{(d)}\cup\{n+1,\ldots,n+d-1\}$. We put these elements into the $|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1$ slots. There are $\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}$ many of places to put the last $d-1$ elements, regardless of their inner order, and exactly one of them is what we needed. After fixed the elements in $n+1,\ldots,n+d-1$, the $|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1$ slots are divided into $d$ parts. We need to put each of $A_{i}^{(d)}$ into the correct place, and the probability of it is
$\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|}^{-1}\binom{|A_{i}^{(2)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(2)}|}^{-1}\cdots\binom{|A_{i}^{(d-1)}|+|A_{i}^{(d)}|}{|A_{i}^{(d-1)}|}^{-1}=\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1},$
and so
$\mathbb{P}(E_{i})=\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}.$
Now it is enough to show $E_{i}$ and $E_{j}$ cannot happen at the same time if $i\neq j$. Actually, if $\sigma\in E_{i}\cap E_{j}$, for every $p<q$, there is a $b\in\{n+1,\ldots,n+d-1\}$ such that $\sigma(a_{i})<\sigma(b)<\sigma(a_{j}),\forall a_{i}\in A_{i}^{(p)},a_{j}\in A_{j}^{(q)}$, so $A_{i}^{(p)}\cap A_{j}^{(q)}=\emptyset$. For the same reason, we have $A_{j}^{(q)}\cap A_{i}^{(p)}=\emptyset$. Then we get a contradiction. Therefore, we have
$1\geqslant\mathbb{P}(\cup_{i=1}^{m}E_{i})=\sum_{i=1}^{m}\mathbb{P}(E_{i})=\sum_{i=1}^{m}\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}.$
Proof of theorem 1.6.
We prove the theorem by induction on $d$. For $d=3$, pick a random permutation $\sigma\in S_{n+1}$ uniformly. $n+1$ will play the role of delimiter. Let $E_{i}$ and $F_{i}$ be events that elements are placed in the order $\sigma(A_{i}^{(1)}),\sigma(n+1),\sigma(A_{i}^{(2)}),\sigma(A_{i}^{(3)})$ and $\sigma(A_{i}^{(1)}),\sigma(A_{i}^{(2)}),\sigma(n+1),\sigma(A_{i}^{(3)})$ correspondingly. In other words, let
$E_{i}$ $=$ $\{\sigma\in S_{n+1}\mid\sigma(a_{1})<\sigma(n+1)<\sigma(a_{2})<\sigma(a_{3}),\forall a_{k}\in A_{i}^{(k)}\},$
$F_{i}$ $=$ $\{\sigma\in S_{n+1}\mid\sigma(a_{1})<\sigma(a_{2})<\sigma(n+1)<\sigma(a_{3}),\forall a_{k}\in A_{i}^{(k)}\}.$
Then similar with the discussion in the former proof, we have
$\mathbb{P}(E_{i})=\mathbb{P}(F_{i})=\left((|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|+1)\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}\right)^{-1}.$
[…]
$\sigma(a_{i})>\sigma(t)>\sigma(a_{j}),\forall a_{i}\in A_{i}^{(3)},a_{j}\in A_{j}^{(2)},$
and $A_{i}^{(3)}\cap A_{j}^{(2)}=\emptyset$. Hence we get a contradiction, and so $E_{i}\cap E_{j}=\emptyset$. Similarly, $F_{i}\cap F_{j}=\emptyset$. Suppose $\sigma\in E_{i}\cap F_{j}$. Then we have $A_{i}^{(2)}\cap A_{j}^{(1)}=A_{i}^{(3)}\cap A_{j}^{(1)}=A_{i}^{(3)}\cap A_{j}^{(2)}=\emptyset$, which is a contradiction. Hence $E_{i}\cap F_{j}=\emptyset$, and similarly $E_{j}\cap F_{i}=\emptyset$.
If $A_{i}^{(2)}\neq\emptyset$, then $E_{i}\cap F_{i}=\emptyset$. Then we have
$1\geqslant\mathbb{P}\left(\bigcup_{A_{i}^{(2)}\neq\emptyset}(E_{i}\cup F_{i})\right)=2\sum_{A_{i}^{(2)}\neq\emptyset}\left((|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|+1)\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}\right)^{-1},$
and so
$\sum_{A_{i}^{(2)}\neq\emptyset}\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}^{-1}\leqslant\frac{n+1}{2}.$
For $A_{i}^{(2)}=\emptyset$, the collection $\{(A_{i}^{(1)},A_{i}^{(3)})\mid A_{i}^{(2)}=\emptyset\}$ forms a Bollobás system (of $d=2$). Then by Bollobás’s Theorem (theorem 1.2), we have
$\sum_{A_{i}^{(2)}=\emptyset}\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}^{-1}=\sum_{A_{i}^{(2)}=\emptyset}\binom{|A_{i}^{(1)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|}^{-1}\leqslant 1.$
Combing these two cases, we have
$\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}^{-1}\leqslant\frac{n+3}{2}.$
Now we are going to prove the theorem for arbitrary $d$ while assuming it is true for the case $d-1$. Pick a random permutation $\sigma\in S_{n+d-2}$ uniformly. Once again, we will use the elements in $\{n+1,\ldots,n+d-2\}$ as delimiters. For $i\in[m]$ and $k\in[d-1]$, let $E_{i}^{(k)}$ be the event that $a_{1}<\cdots<a_{d},\forall a_{l}\in A_{i}^{(l)}$, and all pairs in $\{(A_{i}^{(l)},A_{i}^{(l+1)})\mid l\in[d-1]\}$ except $(A_{i}^{(k)},A_{i}^{(k+1)})$ are divided by a delimiter. Then
$\mathbb{P}(E_{i}^{(k)})=\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-2}{d-2}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}.$
Similar with the discussion in the first part of this proof, we have $E_{i}^{(k)}\cap E_{j}^{(l)}=\emptyset$ for any $i\neq j$, and $E_{i}^{(k)}\cap E_{i}^{(l)}=\emptyset,\forall k\neq l$ if none of $A_{i}^{(2)},\ldots,A_{i}^{(d-1)}$ is empty. Then we have
$\begin{split}1&\geqslant\mathbb{P}\left(\bigcup_{i:A_{i}^{(l)}\neq\emptyset,\forall 2\leqslant l\leqslant d-1}\bigcup_{k=1}^{d-1}E_{i}^{(k)}\right)=\sum_{i:A_{i}^{(l)}\neq\emptyset,\forall 2\leqslant l\leqslant d-1}\sum_{k=1}^{d-1}\mathbb{P}(E_{i}^{(k)})\\ &=(d-1)\sum_{i:A_{i}^{(l)}\neq\emptyset,\forall 2\leqslant l\leqslant d-1}\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-2}{d-2}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1},\end{split}$
and so
$\sum_{i:A_{i}^{(k)}\neq\emptyset,2\leqslant k\leqslant d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}\leqslant\frac{1}{d-1}\binom{n+d-2}{d-2}.$
Note that for $k\in\{2,\ldots,d-1\}$, the family $\{(A_{i}^{(1)},\ldots,A_{i}^{(k-1)},A_{i}^{(k+1)},\ldots,A_{i}^{(d)})\mid A_{i}^{(k)}=\emptyset\}$ is a Bollobás system of $(d-1)$-tuples. Then by the inductive hypothesis, we have
$\sum_{A_{i}^{(k)}=\emptyset}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}\leqslant\frac{1}{d-2}\binom{n+d-3}{d-3}+O(n^{d-4}).$
Hence
$\begin{split}\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}&\leqslant\frac{1}{d-1}\binom{n+d-2}{d-2}+\binom{n+d-3}{d-3}+O(n^{d-4})\\ &=\frac{1}{d-1}\binom{n+d-2}{d-2}+O(n^{d-3}).\end{split}$
3 Proof of theorem 1.9
Proof of theorem 1.9.
Without lose of generality, we assume $A_{i}^{(k)}\subseteq[n],\forall i\in[m],k\in[d]$ for some $n$. Fixed $i\in[m]$, and for any $k\in[d-1]$ let
$f_{ik}(x_{k0},\ldots,x_{ka_{k+1}})=\prod_{t\in\cup_{l=1}^{k}A_{i}^{(l)}}(x_{k0}+tx_{k1}+\cdots+t^{a_{k+1}}x_{ka_{k+1}})$
be an $(a_{k+1}+1)$-variable $(a_{1}+\cdots+a_{k})$-degree homogeneous polynomial. Denote $\underline{x_{k}}=(x_{k0},\ldots,x_{ka_{k+1}})$ for any $k\in[d-1]$, and let
$f_{i}(\underline{x_{1}},\ldots,\underline{x_{d-1}})=f_{i1}(\underline{x_{1}})\cdots f_{i,d-1}(\underline{x_{k-1}}).$
On the other hand, for $i\in[m]$ and $k\in\{2,\ldots,d\}$, let
$B_{ik}=\{(1,t,t^{2},\ldots,t^{a_{k}})\mid t\in A_{i}^{(k)}\}\subseteq\mathbb{R}^{a_{k}+1}.$
Then $\mathrm{span}(B_{ik})$ is a subspace of $\mathbb{R}^{a_{k}+1}$, and $\dim(B_{ik})=a_{k}$. So we can find a nonzero $\underline{\xi_{ik}}\in B_{ik}^{\perp}$. Let $\underline{\xi_{i}}=(\underline{\xi_{i2}},\ldots,\underline{\xi_{id}})$.
Now we claim that
$f_{i}(\xi_{j})\begin{cases}\neq 0,&\textrm{if}~{}i=j;\\ =0,&\textrm{if}~{}i<j.\end{cases}$
Note that $f_{i}(\xi_{j})=f_{i1}(\underline{\xi_{j2}})\cdots f_{i,d-1}(\underline{\xi_{jd}})$, so $f_{i}(\underline{\xi_{j}})=0$ if and only if there is a $k\in[d-1]$ such that $f_{ik}(\underline{\xi_{j,k+1}})=0$. For $i<j$, there exists $p<q$ such that $A_{i}^{(p)}$ and $A_{j}^{(q)}$ are intersecting. Then there is a $t\in A_{i}^{(p)}\cap A_{j}^{(q)}\subseteq\left(A_{i}^{(1)}\cup\cdots\cup A_{i}^{(q-1)}\right)\cap A_{j}^{(q)}$. Let $k=q-1$. By the definition of $\underline{\xi_{j,k+1}}$, the inner product of $(1,t,t^{2},\ldots,t^{a_{k+1}})$ and $\underline{\xi_{j,k+1}}$ is zero, and so $f_{ik}(\underline{\xi_{j,k+1}})=0$. On the other hand, For any $k\in[d-1]$, the fact $A_{i}^{(k+1)}\cap\left(A_{i}^{(1)}\cup\cdots\cup A_{i}^{(k)}\right)=\emptyset$ implies there is not such a $t$, so $f_{ik}(\underline{\xi_{j,k+1}})\neq 0$.
Now we are going to show $f_{1},\ldots,f_{m}$ are linearly independent. Actually, suppose that $c_{1}f_{1}+\cdots+c_{m}f_{m}\equiv 0$ and $c_{1},\ldots,c_{m}$ are not all zeros. Let $k$ be the maximum index such that $c_{m}\neq 0$. Then $f_{m}\equiv-\frac{c_{1}}{c_{m}}f_{1}-\cdots-\frac{c_{m-1}}{c_{m}}f_{m-1}$. But $f_{m}(\underline{\xi_{m}})\neq 0$ and $f_{1}(\underline{\xi_{m}})=\cdots=f_{m-1}(\underline{\xi_{m}})=0$ leads to a contradiction.
Finally, notice that each $f_{ik}$ is an $(a_{k+1}+1)$-variable $(a_{1}+\cdots+a_{k})$-degree homogeneous polynomial, and all of this kind of polynomials (together with zero polynomial) form a $\binom{a_{1}+\cdots+a_{k+1}}{a_{k+1}}$ dimensional vector space, and $f_{i}$ belongs to the tensor product of these spaces. Then the dimension of the full space is
$\binom{a_{1}+a_{2}}{a_{2}}\binom{a_{1}+a_{2}+a_{3}}{a_{3}}\cdots\binom{a_{1}+\cdots+a_{d}}{a_{d}}=\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}},$
and so $m\leqslant\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}$. ∎