### v3

# On Bollobás-type theorems of $d$-tuples

Abstract
In 1965, Bollobás [3] proved that for a Bollobás set-pair system $\{(A_{i},B_{i})\mid i\in[m]\}$, the maximum value of $\sum_{i=1}^{m}\binom{|A_{i}|+|B_{i}|}{A_{i}}^{-1}$ is $1$.  Hegedüs and Frankl [7] recently extended the concept of Bollobás systems to $d$-tuples, conjecturing that for a Bollobás system of $d$-tuples, $\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$, the maximum value of $\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}$ is also $1$. This paper refutes this conjecture and establishes an upper bound for the sum. In the case $d=3$, the derived upper bound is asymptotically tight. Furthermore, we sharpen an inequality for skew Bollobás systems of $d$-tuples in [7], Finally, we determine the maximum size of a uniform skew Bollobás system of $d$-tuples on both sets and spaces.
1 Introduction
Let $[n]=\{1,\ldots,n\}$ be a set of size $n$. The following concepts of Bollobás systems and skew Bollobás systems were introduced by Bollobás [3] and Frankl [4] respectively.
Definition 1.1.
Suppose $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ is a family of pairs of sets, where $A_{i},B_{i}\subseteq[n]$, and $A_{i}\cap B_{i}=\emptyset$. Then $\mathcal{P}$ is called a Bollobás system if $A_{i}\cap B_{j}\neq\emptyset,\forall i\neq j$, and a skew Bollobás system if $A_{i}\cap B_{j}\neq\emptyset,\forall i<j$.
To solve a problem on hypergraphs, Bollobás [3] proved the following theorem in 1965. In 1975, Tarján [19] independently provided a more elegant proof.
Theorem 1.2 (Bollobás [3]).
Let $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ be a Bollobás system, where $A_{i},B_{i}\subseteq[n]$. Then
$\sum_{i=1}^{m}\binom{|A_{i}|+|B_{i}|}{|A_{i}|}^{-1}\leqslant 1.$
Theorem 1.2 also gives the maximum size of a uniform Bollobás system. Specifically, if $|A_{i}|=a,|B_{i}|=b$ for every $i\in[m]$, Theorem 1.2 implies that $m\leqslant\binom{a+b}{a}$. This upper bound was later proved by Jaeger and Payan [8], and Katona [10] independently with different methods. In 1982, using polynomial method, Frankl [4] proved that this upper bound remains true for skew Bollobás systems. This generalization has application in automata theory [16].
Theorem 1.3 (Frankl [4]).
Let $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ be a skew Bollobás system, where $A_{i},B_{i}\subseteq[n]$, and $|A_{i}|=a,|B_{i}|=b$ for any $i$. Then $m\leqslant\binom{a+b}{a}$.
While Theorem 1.2 no longer hold if we replace the condition Bollobás system by skew Bollobás system. The following (tight) inequalities for skew Bollobás systems (instead of Bollobás systems) are proved by Hegedüs and Frankl [7], and Yue [23].
Theorem 1.4 (Hegedüs and Frankl [7], Yue [23]).
Suppose $A_{i},B_{i}\subseteq[n],i\in[m]$, and $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ is a skew Bollobás system. Then we have
$\sum_{i=1}^{m}\left((1+|A_{i}|+|B_{i}|)\binom{|A_{i}|+|B_{i}|}{|A_{i}|}\right)^{-1}\leqslant 1,$
and so
$\sum_{i=1}^{m}\binom{|A_{i}|+|B_{i}|}{|A_{i}|}^{-1}\leqslant 1+n.$
Note that there are two kinds of structures: Bollobás systems and skew Bollobás systems; and for each of them, there are two questions to ask: what is the maximum size of the family in the uniform case, and what inequality can we get in the nonuniform case. The following table is a brief summary so far.
Bollobás systems Skew Bollobás systems
Uniform cases Bollobás [3] Hegedüs and Frankl [7] and Yue [23]
Nonuniform cases Bollobás [3] Frankl [4]
Table 1: A brief map for Bollobás-type theorem
A natural generalization of (skew) Bollobás systems involves considering families of $d$-tuples instead of set-pairs. In [7], Hegedüs and Frankl defined Bollobás systems and skew Bollobás systems for $d$-tuples as follows. Note that there are some other ways to define these concepts on $d$-tuples, see [15] for instance.
Definition 1.5.
Let =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ be a collection of $d$-tuples. Suppose that $A_{i}^{(p)}\cap A_{i}^{(q)}=\emptyset$ for any $i\in[m]$ and $p\neq q$. Then $\mathcal{F}$ is a Bollobás system (skew Bollobás system) if for any $i\neq j$ ($i<j$), there exist $p<q$ such that $A_{i}^{(p)}\cap A_{j}^{(q)}\neq\emptyset$.
For a $d$-tuple $(A^{(1)},\ldots,A^{(d)})$, we say $(|A^{(1)}|,\ldots,|A^{(d)}|)$ is its type. Following is an example of Bollobás system of $d$-tuples.
Example 1.
For fixed $a_{1},\ldots,a_{d}$, let $\mathcal{F}$ be the collection of all disjoint $d$-tuples $(A^{(1)},\ldots,A^{(d)})$ of type $(a_{1},\ldots,a_{d})$ such that $A^{(k)}\in[a_{1}+\cdots+a_{d}],\forall k\in[d]$. For distinct $(A^{(1)},\ldots,A^{(d)})$ and $(B^{(1)},\ldots,B^{(d)})$, there is a $k$ such that $A^{(k)}\neq B^{(k)}$. Then we have $A^{(k)}\cap\left(\cup_{l\neq k}B^{(l)}\right)\neq\emptyset$ and $B^{(k)}\cap\left(\cup_{l\neq k}A^{(l)}\right)\neq\emptyset$. Hence $\mathcal{F}$ is a Bollobás system of $d$-tuples.
For a fixed $n\geqslant 0$, and $k_{1},\ldots,k_{t}\leqslant n$, define
$\binom{n}{k_{1},\ldots,k_{t}}:=\frac{n!}{k_{1}!\cdots k_{t}!(n-k_{1}-\cdots-k_{t})!},$
which is a natural generalization of the binomial coefficients. Then the size of the Bollobás system defined in Example 1 is $\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}$.
To generalize Theorem 1.2, Hegedüs and Frankl [7] proposed the following conjecture.
Conjecture 1 (Hegedüs and Frankl [7]).
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a Bollobás system of $d$-tuples. Then
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}}^{-1}\leqslant 1.$
However, we can deny this conjecture with the following example.
Example 2.
Fix an arbitrary $n$, and $d=3$. For $l=0,\dots,\lfloor\frac{n}{2}\rfloor$, let $\mathcal{F}_{l}$ be the collection of all disjoint triples $(A,B,C)$ of type $(l,n-2l,l)$, such that $A,B,C\subseteq[n]$, and $\mathcal{F}$ be the union of $\mathcal{F}_{l}$’s. We claim that $\mathcal{F}$ is a Bollobás system. Actually, for two different families $(A,B,C),(A^{\prime},B^{\prime},C^{\prime})\in\mathcal{F}_{l}$, if $A=A^{\prime}$, then $B\cap C^{\prime}\neq\emptyset$ and $B^{\prime}\cap C\neq\emptyset$. If $A\neq A^{\prime}$, then $A\cap(B^{\prime}\cup C^{\prime})=\neq\emptyset$ and $A^{\prime}\cap(B\cup C)\neq\emptyset$. For $(A,B,C)\in\mathcal{F}_{l}$ and $(A^{\prime},B^{\prime},C^{\prime})\in\mathcal{F}_{l^{\prime}}$ where $l<l^{\prime}$, note that $|A|+|B|+|C^{\prime}|=|A^{\prime}|+|B|+|C|=n-l+l^{\prime}>n$, so $A\cap(B^{\prime}\cup C^{\prime})\neq\emptyset$ and $A^{\prime}\cap(B\cup C)\neq\emptyset$. Hence $\mathcal{F}$ is a Bollobás system. But we have
$\sum_{(A,B,C)\in\mathcal{F}}{\binom{|A|+|B|+|C|}{|A|,|B|,|C|}}^{-1}=\sum_{l=0}^{\lfloor\frac{n}{2}\rfloor}\sum_{(A,B,C)\in\mathcal{F}_{l}}{\binom{|A|+|B|+|C|}{|A|,|B|,|C|}}^{-1}=\left\lfloor\frac{n}{2}\right\rfloor+1.$
It is important to note that Tian and Wu [20] also notice the invalidity of this conjecture, as well as Theorem 1.8. For further details, please refer to the remark at the end of this paper.
Since Conjecture 1 is not true, a natural subsequent question arises: for a Bollobás system $\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ of $d$-tuples, what is the maximum value of $\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}$? Generalizing the probabilistic method in [23], we can prove the following inequalities for Bollobás systems of $d$-tuples. The inequality for $d=3$ shows the construction in Example 2 is nearly extremum, and is asymptotically tight as $n\rightarrow\infty$. For the case $d>3$, it is still not clear if the inequality is tight or not.
Theorem 1.6.
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)}\mid i\in[m])\}$ is a Bollobás system of $d$-tuples, and $A_{i}^{(p)}\subseteq[n]$ for all $i\in[m],p\in[d]$. Then for the case $d=3$ we have
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}}^{-1}\leqslant\frac{n+3}{2},$
and for arbitrary $d$ we have
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}}^{-1}\leqslant\frac{1}{d-1}\binom{n+d-2}{d-2}+O(n^{d-3})$
as $n\rightarrow\infty$.
For skew Bollobás systems of $d$-tuples, Hegedüs and Frankl [7] proved the following theorem.
Theorem 1.7 (Hegedüs and Frankl [7]).
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a skew Bollobás system of $d$-tuples, and $A_{i}^{(p)}\subseteq[n]$ for all $i\in[m]$ and $p\in[d]$. Then
$\sum_{i=1}^{m}{\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}}^{-1}\leqslant\binom{n+d-1}{d-1}.$
Although this result is already tight, using similar probabilistic argument, we can still improve it as follows.
Theorem 1.8.
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a skew Bollobás system of $d$-tuples. Then
$\sum_{i=1}^{m}\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}\leqslant 1.$
Note that $|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|\leqslant n$, so Theorem 1.8 implies Theorem 1.7. Theorem 1.8 also provides an upper bound for the size of uniform skew Bollobás systems. If we further assume that $|A_{i}^{(p)}|=a_{p},\forall i\in[m],p\in[d]$, then we have $m\leqslant\binom{n+d-1}{d-1}\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}$ where $n$ denotes the size of the ground set. However, this upper bound is not so good. Actually, we can prove the following upper bound, which is confirmed to be tight by Example 1. (Note that a Bollobás system is also a skew Bollobás system.)
Theorem 1.9.
Suppose that =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ is a skew Bollobás system of $d$-tuples, and $|A_{i}^{(p)}|=a_{p},\forall i\in[m],p\in[d]$. Then
$m\leqslant\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}.$
Like Table 1, the following table give a brief summary for Bollobás-type theorem on $d$-tuples.
Bollobás systems Skew Bollobás systems
Uniform Solved: Theorem 1.9 Solved: Theorem 1.9
Nonuniform Conjecture: Hegedüs and Frankl [7] Solved: Hegedüs and Frankl [7]
Partly solved: Example 2 and Theorem 1.6 Improved: Theorem 1.8
$\star$: Tian and Wu [20] also have contribution on these two questions.
Table 2: A brief map for Bollobás-type theorem on $d$-tuples
As an analog, Bollobás systems and skew Bollobás systems of vector spaces are defined as follows.
Definition 1.10.
Let $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ be a family of pairs of subspaces of a fixed vector space $V$, where $\dim(A_{i}\cap B_{i})=0$. We say that $\mathcal{P}$ is a Bollobás system if $\dim(A_{i}\cap B_{j})>0,\forall i\neq j$, and a skew Bollobás system if $\dim(A_{i}\cap B_{j})>0,\forall i<j$.
In 1977, employing powerful algebraic techniques, Lovász [12, 13, 14] extended Theorem 1.3 to the domain of matroids. This seminal extension not only opened a new avenue for solving combinatorial problems using the exterior product method, but also elevated the Bollobás-type theorem to a central position within the field of extremal set theory. Perhaps for the sake of simplicity, Füredi [5] reformulated it as follows for real vector spaces.
Theorem 1.11 (Lovász [12, 13, 14]).
Let $\mathcal{P}=\{(A_{i},B_{i})\mid i\in[m]\}$ be a (skew) Bollobás system of subspaces of $\mathbb{R}^{n}$. If $\dim(A_{i})=a,\dim(B_{i})=b$ for all $i$, then $m\leqslant\binom{a+b}{a}$.
The conclusion of Theorem 1.11 holds for vector spaces over an arbitrary field. For the case of $\mathbb{Q}^{n}$ or $\mathbb{C}^{n}$, the proofs remain essentially the same. For other fields, such as finite fields, some technical adaptations may be necessary. More details about this can be found in [2]. After Füredi, other scholars (such as Scott and Wilmer [17], and Yu, Kong, Xi, Zhang, and Ge [22]) have tended to express their theorems in the same way, even though they may hold more generally.
Bollobás-type theorems and their variations have attracted significant interest among mathematicians. In 1984, Füredi [5] proved the threshold (or $t$-intersecting) version of Bollobás-type theorem (for both sets and spaces). This work laid the foundation for further extensions by Zhu [24], Talbot [18], and Kang, Kim, and Kim [9]. Recent research has explored even more diverse variations of Bollobás-type theorems, such as those concerning partitions of sets [1], affine spaces [6], and weakly Bollobás systems [11, 21].
It is natural to extend the concept of Bollobás systems and skew Bollobás systems to $d$-tuples of spaces.
Definition 1.12.
Let =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$, where $A_{i}^{(k)}\leqslant V=\mathbb{R}^{n}$. Suppose that $\dim(A_{i}^{(1)}+\cdots+A_{i}^{(d)})=\dim(A_{i}^{(1)})+\cdots+\dim(A_{i}^{(d)})$ for any $i\in[m]$. Then $\mathcal{F}$ is a Bollobás system (skew Bollobás system) if for any $i\neq j$ ($i<j$), there exist $p<q$ such that $A_{i}^{(p)}\cap A_{j}^{(q)}\neq\{0\}$.
Our last result is the following theorem, which determines the maximum size of a skew Bollobás system of $d$-tuples of spaces. This theorem strengthens Theorem 1.9, and generalizes Theorem 1.11.
Theorem 1.13.
Let =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$ be a skew Bollobás system of $d$-tuples of spaces, where $A_{i}^{(k)}\leqslant V=\mathbb{R}^{n}$. Suppose that $\dim(A_{i}^{(k)})=a_{k}$ for every $i\in[m]$ and $k\in[d]$. Then we have
$m\leqslant\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}.$
This paper is organized as follows: in Section 2, we prove Theorem 1.6 and Theorem 1.8, in Section 3, we prove Theorem 1.13, and obtain Theorem 1.9 from it.
2 Proof of Theorem 1.6 and Theorem 1.8
For a fixed $n>0$, denote $S_{n}$ the group of all permutations on $[n]$. For a $\sigma\in S_{n}$ and $A\subseteq[n]$, denote $\sigma(A):=\{\sigma(a)\mid a\in A\}$ the image of $A$ over $\sigma$. For $\sigma\in S_{n}$ and $A,B\subseteq[n]$, obviously $\sigma(A)\cap\sigma(B)=\emptyset$ if and only if $A\cap B=\emptyset$. Further more, for a family of $d$-tuples =$\mathcal{F}=\{(A_{i}^{(1)},\ldots,A_{i}^{(d)})\mid i\in[m]\}$, where $A_{i}^{(k)}\subseteq[n],\forall i\in[m],k\in[d]$, and any $\sigma\in S_{n}$,
=$\mathcal{F}^{\prime}=\{(\sigma(A_{i}^{(1)}),\ldots,\sigma(A_{i}^{(d)}))\mid i\in[m]\}$
is a (skew) Bollobás system, if and only if $\mathcal{F}$ is a (skew) Bollobás system. We will prove Theorem 1.6 and Theorem 1.8 with the tool of random permutations, and begin with Theorem 1.8.
Proof of Theorem 1.8.
Without lose of generality, we assume $A_{i}^{(k)}\subseteq[n],\forall i\in[m],k\in[d]$ for some $n$. Pick a random permutation $\sigma\in S_{n+d-1}$ uniformly, and it gives the $n+d-1$ elements a new order. We will use the elements in $\{n+1,\ldots,n+d-1\}$ as delimiters. Let $E_{i}$ be the event that elements in $\sigma(A_{i}^{(1)}),\ldots,\sigma(A_{i}^{(d)})$ appear in order, and each pair of neighbors are divided by a delimiter. Formally speaking, define
$\begin{split}E_{i}:=&\{\sigma\in S_{n+d-1}\mid\sigma(a_{1})<\sigma(b_{1})<\sigma(a_{2})<\cdots<\sigma(b_{d-1})<\sigma(a_{d}),\forall a_{k}\in A_{i}^{(k)},\\ &\textrm{where}~{}\{b_{1},\ldots,b_{d-1}\}=\{n+1,\ldots,n+d-1\}\}.\end{split}$
To calculate the probability of $E_{i}$, we only need to consider about elements in $A_{i}^{(1)}\cup\cdots\cup A_{i}^{(d)}\cup\{n+1,\ldots,n+d-1\}$. We put these elements into the $|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1$ slots. There are $\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}$ many of places to put the last $d-1$ elements, regardless of their inner order, and exactly one of them is what we needed. After fixed the elements in $\{n+1,\ldots,n+d-1\}$, the rest $|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|$ slots are divided into $d$ parts. We need to put each of $A_{i}^{(k)}$ into the correct place, and the probability of it is
$\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|}^{-1}\binom{|A_{i}^{(2)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(2)}|}^{-1}\cdots\binom{|A_{i}^{(d-1)}|+|A_{i}^{(d)}|}{|A_{i}^{(d-1)}|}^{-1}=\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1},$
and so
$\mathbb{P}(E_{i})=\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}.$
Now it is enough to show $E_{i}$ and $E_{j}$ cannot happen at the same time if $i\neq j$. Actually, if $\sigma\in E_{i}\cap E_{j}$, for every $p<q$, there is a $b\in\{n+1,\ldots,n+d-1\}$ such that $\sigma(a_{i})<\sigma(b)<\sigma(a_{j}),\forall a_{i}\in A_{i}^{(p)},a_{j}\in A_{j}^{(q)}$, so $A_{i}^{(p)}\cap A_{j}^{(q)}=\emptyset$. For the same reason, we have $A_{i}^{(q)}\cap A_{j}^{(p)}=\emptyset$. Then we get a contradiction, no matter $i<j$ or $i>j$. Therefore, we have
$1\geqslant\mathbb{P}\left(\bigcup_{i=1}^{m}E_{i}\right)=\sum_{i=1}^{m}\mathbb{P}(E_{i})=\sum_{i=1}^{m}\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-1}{d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}.$
Now we are going to prove Theorem 1.6 with similar probabilistic argument.
Proof of Theorem 1.6.
We prove the theorem by induction on $d$. For $d=3$, pick a random permutation $\sigma\in S_{n+1}$ uniformly. $n+1$ will play the role of delimiter. Let $E_{i}$ and $F_{i}$ be events that elements are placed in the order $\sigma(A_{i}^{(1)}),\sigma(n+1),\sigma(A_{i}^{(2)}),\sigma(A_{i}^{(3)})$ and $\sigma(A_{i}^{(1)}),\sigma(A_{i}^{(2)}),\sigma(n+1),\sigma(A_{i}^{(3)})$ correspondingly. In other words, let
$E_{i}$ ::=:= $\{\sigma\in S_{n+1}\mid\sigma(a_{1})<\sigma(n+1)<\sigma(a_{2})<\sigma(a_{3}),\forall a_{k}\in A_{i}^{(k)}\},$
$F_{i}$ ::=:= $\{\sigma\in S_{n+1}\mid\sigma(a_{1})<\sigma(a_{2})<\sigma(n+1)<\sigma(a_{3}),\forall a_{k}\in A_{i}^{(k)}\}.$
Then similar with the discussion in the former proof, we have
$\mathbb{P}(E_{i})=\mathbb{P}(F_{i})=\left((|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|+1)\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}\right)^{-1}.$
[…]
$\sigma(a_{i})>\sigma(t)>\sigma(a_{j}),\forall a_{i}\in A_{i}^{(3)},a_{j}\in A_{j}^{(2)},$
and $A_{i}^{(3)}\cap A_{j}^{(2)}=\emptyset$. Hence we get a contradiction, and so $E_{i}\cap E_{j}=\emptyset$. Similarly, $F_{i}\cap F_{j}=\emptyset$. Suppose $\sigma\in E_{i}\cap F_{j}$. Then elements in $\sigma(A_{i}^{(1)}),\sigma(A_{j}^{(1)}),\sigma(A_{j}^{(2)})$ appear before $\sigma(n+1)$, while that in $\sigma(A_{i}^{(2)}),\sigma(A_{i}^{(3)}),\sigma(A_{j}^{(3)})$ after it, so we have $A_{i}^{(2)}\cap A_{j}^{(1)}=A_{i}^{(3)}\cap A_{j}^{(1)}=A_{i}^{(3)}\cap A_{j}^{(2)}=\emptyset$, which is a contradiction. Hence $E_{i}\cap F_{j}=\emptyset$, and similarly $E_{j}\cap F_{i}=\emptyset$.
If $A_{i}^{(2)}\neq\emptyset$, then $E_{i}\cap F_{i}=\emptyset$. Then we have
$1\geqslant\mathbb{P}\left(\bigcup_{A_{i}^{(2)}\neq\emptyset}(E_{i}\cup F_{i})\right)=2\sum_{A_{i}^{(2)}\neq\emptyset}\left((|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|+1)\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}\right)^{-1},$
and so
$\sum_{A_{i}^{(2)}\neq\emptyset}\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}^{-1}\leqslant\frac{n+1}{2}.$
For $A_{i}^{(2)}=\emptyset$, the collection $\{(A_{i}^{(1)},A_{i}^{(3)})\mid A_{i}^{(2)}=\emptyset\}$ forms a Bollobás system (of $d=2$). Then by Theorem 1.2, we have
$\sum_{A_{i}^{(2)}=\emptyset}\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}^{-1}=\sum_{A_{i}^{(2)}=\emptyset}\binom{|A_{i}^{(1)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|}^{-1}\leqslant 1.$
Combing these two cases, we have
$\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+|A_{i}^{(2)}|+|A_{i}^{(3)}|}{|A_{i}^{(1)}|,|A_{i}^{(2)}|,|A_{i}^{(3)}|}^{-1}\leqslant\frac{n+3}{2}.$
Now we are going to prove the theorem for arbitrary $d$ while assuming it is true for the case $d-1$. Pick a random permutation $\sigma\in S_{n+d-2}$ uniformly. Once again, we will use the elements in $\{n+1,\ldots,n+d-2\}$ as delimiters. For $i\in[m]$ and $k\in[d-1]$, let $E_{i}^{(k)}$ be the event that $\sigma(a_{1})<\cdots<\sigma(a_{d}),\forall a_{l}\in A_{i}^{(l)}$, and all pairs in $\{(\sigma(A_{i}^{(l)}),\sigma(A_{i}^{(l+1)}))\mid l\in[d-1]\}$ except $(\sigma(A_{i}^{(k)}),\sigma(A_{i}^{(k+1)}))$ are divided by a delimiter. Then
$\mathbb{P}(E_{i}^{(k)})=\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-2}{d-2}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1}.$
Similar with the discussion in the first part of this proof, we have $E_{i}^{(k)}\cap E_{j}^{(l)}=\emptyset$ for any $i\neq j$, and $E_{i}^{(k)}\cap E_{i}^{(l)}=\emptyset,\forall k\neq l$ if none of $A_{i}^{(2)},\ldots,A_{i}^{(d-1)}$ is empty. Then we have
$\begin{split}1&\geqslant\mathbb{P}\left(\bigcup_{i:A_{i}^{(l)}\neq\emptyset,\forall 2\leqslant l\leqslant d-1}\bigcup_{k=1}^{d-1}E_{i}^{(k)}\right)=\sum_{i:A_{i}^{(l)}\neq\emptyset,\forall 2\leqslant l\leqslant d-1}\sum_{k=1}^{d-1}\mathbb{P}(E_{i}^{(k)})\\ &=(d-1)\sum_{i:A_{i}^{(l)}\neq\emptyset,\forall 2\leqslant l\leqslant d-1}\left(\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|+d-2}{d-2}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}\right)^{-1},\end{split}$
and so
$\sum_{i:A_{i}^{(k)}\neq\emptyset,2\leqslant k\leqslant d-1}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}\leqslant\frac{1}{d-1}\binom{n+d-2}{d-2}.$
Note that for $k\in\{2,\ldots,d-1\}$, the family $\{(A_{i}^{(1)},\ldots,A_{i}^{(k-1)},A_{i}^{(k+1)},\ldots,A_{i}^{(d)})\mid A_{i}^{(k)}=\emptyset\}$ is a Bollobás system of $(d-1)$-tuples. Then by the inductive hypothesis, we have
$\sum_{A_{i}^{(k)}=\emptyset}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}\leqslant\frac{1}{d-2}\binom{n+d-3}{d-3}+O(n^{d-4}).$
Hence
$\begin{split}\sum_{i=1}^{m}\binom{|A_{i}^{(1)}|+\cdots+|A_{i}^{(d)}|}{|A_{i}^{(1)}|,\ldots,|A_{i}^{(d)}|}^{-1}&\leqslant\frac{1}{d-1}\binom{n+d-2}{d-2}+\binom{n+d-3}{d-3}+O(n^{d-4})\\ &=\frac{1}{d-1}\binom{n+d-2}{d-2}+O(n^{d-3}).\end{split}$
3 Proof of Theorem 1.9 and Theorem 1.13
A handy tool to deal with (skew) Bollobás systems of spaces is exterior algebra. Suppose that $\alpha_{1},\ldots,\alpha_{k}$ are $n$-dimensional (raw) vectors in $\mathbb{R}^{n}$, let
$A=\begin{pmatrix}\alpha_{1}\\ \vdots\\ \alpha_{k}\end{pmatrix}$
be a $k\times n$ matrix. For $I\in\binom{[n]}{k}$, denote $A_{I}$ the $k\times k$ submatrix of $A$ consisting of the $k$ columns labeled by $I$. Then the exterior product (or wedge product) of $\alpha_{1},\ldots,\alpha_{k}$ is an $\binom{n}{k}$-dimensional vector
$\alpha_{1}\wedge\cdots\wedge\alpha_{k}:=(|A_{I}|\mid I\in\binom{[n]}{k}).$
Note that if $n=k$ then the exterior product is just the determinant of $A$.
The above simplified version of the definition of exterior product comes from [2], in where more details and propositions can be found. For the original abstract definition of exterior product and exterior algebra, one may refer to any textbooks of multilinear algebra. Here we emphasize the following two lemmas, which are basic conclusions in multilinear algebra, but play a vital role in our proof.
Lemma 3.1.
Let $V=\mathbb{R}^{n}$, and $V^{k}$ be the Cartesian product of $k$ many of $V$. Then
$f:V^{k}\rightarrow\mathbb{R}^{\binom{n}{k}},\quad(\alpha_{1},\ldots,\alpha_{k})\mapsto\alpha_{1}\wedge\cdots\wedge\alpha_{k}$
is a $k$-linear mapping.
Lemma 3.2.
Suppose that $\alpha_{1},\ldots,\alpha_{k}$ are $k$ many of $n$-dimensional vectors. Then they are linearly independent if and only if $\alpha_{1}\wedge\cdots\wedge\alpha_{k}\neq 0$.
Suppose $\alpha=\alpha_{1}\wedge\cdots\wedge\alpha_{s}$ and $\beta=\beta_{1}\wedge\cdots\wedge\beta_{t}$, denote
$\alpha\wedge\beta:=\alpha_{1}\wedge\cdots\wedge\alpha_{s}\wedge\beta_{1}\wedge\cdots\wedge\beta_{t},$
which is a bilinear mapping.
Fix $V=\mathbb{R}^{n}$, and a $k\leqslant n$. Denote $\bigwedge^{k}V$, the collection of the exterior products of $k$ many of vectors in $V$. Then $\bigwedge^{k}V$ is a $\binom{n}{k}$-dimensional vector space. Let $W$ be a $k$-dimensional subspace of $V$, and $\alpha_{1},\ldots,\alpha_{k}$ be a basis of $W$. Denote
$\bigwedge W:=\alpha_{1}\wedge\cdots\wedge\alpha_{k}\in\bigwedge^{k}V.$
Note that this notation is not well defined, for $\bigwedge W$ depends on the choice of the basis. However, the only difference between them is a nonzero factor. So we can use it harmlessly as long as we only care about if the result is zero or not.
To prove Theorem 1.9 and Theorem 1.13, we also need the following concept and lemma for general position argument. Suppose $V$ and $W$ are $n$-dimensional and $k$-dimensional vector spaces over field $F$, and $U_{1},\cdots,U_{m}$ are subspaces of $V$, where $\dim(U_{i})=r_{i}$. We say linear mapping $\phi:V\rightarrow W$ is in general position with $U_{i}$’s, if
$\dim(\phi(U_{i}))=\min\{r_{i},k\},\quad i=1,\ldots,m.$
Note that for every subspace $U_{i}$$\min\{r_{i},k\}$ is the maximum possible dimension of $\phi(U_{i})$. The following lemma shows that for fixed $V$$W$, and $U_{i}$’s, a linear mapping $\phi:V\rightarrow W$ that in general position with $U_{i}$’s always exists, as long as $|F|$ is large enough.
Lemma 3.3 ([2]).
Suppose $V$ and $W$ are $n$-dimensional and $k$-dimensional vector spaces over field $F$$U_{1},\ldots,U_{m}$ are subspaces of $V$, and $\dim(U_{i})=r_{i}$. If $|F|>(n-k)(m+1)$ then there exists a linear mapping $\phi:V\rightarrow W$ that is in general position with $U_{i}$’s.
Now we are ready to prove Theorem 1.13
Proof of Theorem 1.13.
For every $k\in[d]$, let $V_{k}=\mathbb{R}^{a_{1}+\cdots+a_{k}}$. By general position argument (Lemma 3.3), there exists a linear mapping $\phi_{k}:V\rightarrow V_{k}$ that maintains the dimension of $A_{i}^{(1)}+\cdots+A_{i}^{(k)}$ for $i\in[m]$, and the dimension of $A_{i}^{(p)}+A_{j}^{(q)}$ for $i,j\in[m]$ and $p,q\in[k]$. The first condition leads to the fact that $V_{k}=\phi_{k}(A_{i}^{(1)})\bigoplus\cdots\bigoplus\phi_{k}(A_{i}^{(k)})$, and the second condition guarantees that
$\dim(\phi_{k}(A_{i}^{(p)})\cap\phi_{k}(A_{j}^{(q)}))=\dim(A_{i}^{(p)}\cap A_{j}^{(q)}),\quad\forall i,j\in[m],p,q\in[k].$ (1)
Actually, we have
$\begin{split}\dim(\phi_{k}(A_{i}^{(p)})\cap\phi_{k}(A_{j}^{(q)}))&=\dim(\phi_{k}(A_{i}^{(p)}))+\dim(\phi_{k}(A_{j}^{(q)}))-\dim(\phi_{k}(A_{i}^{(p)})+\phi_{k}(A_{j}^{(q)}))\\ &=\dim(A_{i}^{(p)})+\dim(A_{j}^{(q)})-\dim(\phi_{k}(A_{i}^{(p)}+A_{j}^{(q)}))\\ &=\dim(A_{i}^{(p)})+\dim(A_{j}^{(q)})-\dim(A_{i}^{(p)}+A_{j}^{(q)})\\ &=\dim(A_{i}^{(p)}\cap A_{j}^{(q)}).\end{split}$
For any $i\in[m],k\in[d],p\in[k]$$\phi_{k}(A_{i}^{(p)})$ is a subspace of $V_{k}$. Let $\alpha(i,k,p)=\bigwedge\phi_{k}(A_{i}^{(p)})\in\bigwedge^{a_{p}}V_{k}$. For $i\in[m]$ and $k\in\{2,\ldots,d\}$, let
$f_{i,k}:\bigwedge^{a_{k}}V_{k}\rightarrow\mathbb{R},\quad\beta_{k}\mapsto\left(\alpha(i,k,1)\wedge\alpha(i,k,2)\wedge\cdots\wedge\alpha(i,k,k-1)\right)\wedge\beta_{k}$
be a linear mapping. Finally, for every $i\in[m]$, let
$f_{i}(\beta_{2},\ldots,\beta_{d}):=f_{i,2}(\beta_{2})\cdots f_{i,d}(\beta_{d}),$
and
$\xi_{i}:=(\alpha(i,2,2),\alpha(i,3,3),\ldots,\alpha(i,d,d))\in\bigwedge^{a_{2}}V_{2}\times\cdots\times\bigwedge^{a_{d}}V_{d}.$
Now we claim that $f_{i}(\xi_{i})\neq 0$ and $f_{i}(\xi_{j})=0,\forall i<j$. Actually, for every $k\in\{2,\ldots,d\}$, we have $V_{k}=\phi_{k}(A_{i}^{(1)})\bigoplus\cdots\bigoplus\phi_{k}(A_{i}^{(k)})$. Then we can form a basis of $V_{k}$ by putting together the bases of $\phi_{k}(A_{i}^{(p)})$ for every $p\in[k]$. Hence
$f_{i,k}(\alpha(i,k,k))=\alpha(i,k,1)\wedge\alpha(i,k,2)\wedge\cdots\wedge\alpha(i,k,k-1)\wedge\alpha(i,k,k)$
is nonzero, and so $f_{i}(\xi_{i})\neq 0$. On the other hand, for every $i<j$, there exist $p<q$ such that
$\dim(\phi_{q}(A_{i}^{(p)})\cap\phi_{q}(A_{j}^{(q)}))=\dim(A_{i}^{(p)}\cap A_{j}^{(q)})>0.$
Then we have $\alpha(i,q,p)\wedge\alpha(j,q,q)=0$, so
$f_{i,q}(\alpha(i,q,q))=(\alpha(i,q,1)\wedge\alpha(i,q,2)\wedge\cdots\wedge\alpha(i,q,q-1))\wedge\alpha(j,q,q)=0,$
and $f_{i}(\xi_{j})=0$.
This is enough to show that $f_{1},\ldots,f_{m}$ are linearly independent. Actually, suppose that $c_{1}f_{1}+\cdots+c_{m}f_{m}\equiv 0$ and $c_{1},\ldots,c_{m}$ are not all zeros. Let $k$ be the maximum index such that $c_{m}\neq 0$. Then $f_{m}\equiv-\frac{c_{1}}{c_{m}}f_{1}-\cdots-\frac{c_{m-1}}{c_{m}}f_{m-1}$. But $f_{m}(\xi_{m})\neq 0$ and $f_{1}(\xi_{m})=\cdots=f_{m-1}(\xi_{m})=0$ leads to a contradiction.
Note that for fixed $i\in[m]$, and each $k\in\{2,\ldots,d\}$$f_{i,k}$ is a linear mapping from a $\binom{a_{1}+\cdots+a_{k}}{a_{k}}$ dimensional space to $\mathbb{R}$, and all of this kind of mapping forms a $\binom{a_{1}+\cdots+a_{k}}{a_{k}}$ dimensional space. In addition, $f_{i}$ belongs to the tensor product of these spaces, so we have
$m\leqslant\binom{a_{1}+a_{2}}{a_{2}}\binom{a_{1}+a_{2}+a_{3}}{a_{3}}\cdots\binom{a_{1}+\cdots+a_{d}}{a_{d}}=\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}.$
Finally, we show how can we obtain Theorem 1.9 from Theorem 1.13.
Proof of Theorem 1.9.
Without lose of generality, we may assume $A_{i}^{k}\subseteq[n],\forall i\in[m],k\in[d]$ for some $n$. Let $\epsilon_{1},\ldots,\epsilon_{n}$ be a basis of $\mathbb{R}^{n}$, and for every set $A\subseteq[n]$, let $s(A)=\mathrm{span}(\{\epsilon_{a}\mid a\in A\})$ be a subspace of $\mathbb{R}^{n}$. Consider the family
:=$\mathcal{F}^{\prime}:=\{(s(A_{i}^{(1)}),\ldots,s(A_{i}^{(d)}))\mid i\in[m]\},$
we have $\dim(s(A_{i}^{(k)}))=|A_{i}^{(k)}|=a_{k},\forall i\in[m],k\in[d]$, and $\dim(s(A_{i}^{(p)})\cap s(A_{j}^{(q)}))=|A_{i}^{(p)}\cap A_{j}^{(q)}|,\forall i,j\in[m],p,q\in[d]$. Then the fact that $\mathcal{F}$ is a skew Bollobás system (of $d$-tuples of sets) guarantees that $\mathcal{F}^{\prime}$ is a skew Bollobás system (of $d$-tuples of spaces). Hence
$|\mathcal{F}|=|\mathcal{F}^{\prime}|\leqslant\binom{a_{1}+\cdots+a_{d}}{a_{1},\ldots,a_{d}}.$
Remark
It was several months ago that the study in this paper began, and the counter example of Conjecture 1 (Example 2) was found. The first version of this paper was uploaded to arXiv on 2024-11-26. In the study of Tian and Wu [20], they proved the invalidness of Conjecture 1 using Young’s lattice. They didn’t put their manuscript to arXiv so we didn’t notice their work, until Prof. Katona forwarded us their slide used in a conference. With the help of Young’s lattice, they went further on the construction, and they guessed their construction is the maximum case, but they were not able to prove an upper bound in the general case. Theorem 1.6 in this paper gave an upper bound, and partly answered their question. Although there is still a gap between the upper bound in Theorem 1.6 and the construction in [20], the combination of these two research shows that their construction (even if not extremum) and our upper bound (even if not tight) are not so bad. In addition, as I know, both of us discovered Theorem 1.8 independently.
Acknowledgments
We would like to show our appreciate to Prof. Gyula O. H. Katona for let us know the study of Tian and Wu.